### v2

# USTC-TD: A Test Dataset and Benchmark for Image and Video Coding in 2020s

Abstract
Image/video coding has been a remarkable research area for both academia and industry for many years. Testing datasets, especially high-quality image/video datasets are desirable for the justified evaluation of coding-related research, practical applications, and standardization activities. We put forward a test dataset namely USTC-TD, which has been successfully adopted in the practical end-to-end image/video coding challenge of the IEEE International Conference on Visual Communications and lmage Processing (VCIP) in 2022 and 2023. USTC-TD contains 40 images at 4K spatial resolution and 10 video sequences at 1080p spatial resolution, featuring various content due to the diverse environmental factors (e.g. scene type, texture, motion, view) and the designed imaging factors (e.g. illumination, lens, shadow). We quantitatively evaluate USTC-TD on different image/video features (spatial, temporal, color, lightness), and compare it with the previous image/video test datasets, which verifies the wider coverage and more diversity of the proposed dataset. We also evaluate both classic standardized and recent learned image/video coding schemes on USTC-TD with PSNR and MS-SSIM, and provide an extensive benchmark for the evaluated schemes. Based on the characteristics and specific design of the proposed test dataset, we analyze the benchmark performance and shed light on the future research and development of image/video coding. All the data are released online: https://esakak.github.io/USTC-TD .
Index Terms:
Benchmark, image coding, standardization, test dataset, video coding.
I Introduction
Nowadays, with the dramatic growth of data traffic over the internet and the emergent application of versatile image/video formats such as 2K, 4K, high dynamic range, and wide color gamut, there is a pressing demand for storage and transmission. To address this challenge, in recent decades, image/video compression is employed to reduce the amount of data significantly, and several video coding standards have been developed, such as High Efficiency Video Coding (H.265/HEVC)[1], Versatile Video Coding (H.266/VVC)[2], Audio Video Standard (AVS1, AVS2, AVS3)[3], AOMedia Video 1 and 2 (AV1[4], AV2[5]).
End-to-end image/video compression has been a research focus on visual data compression for both academia and industry for over six years[6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. A number of technologies have been developed, such as expressive auto-encoder neural networks, precise probability estimation neural networks, and conditional end-to-end coding frameworks, and so on. Until recently, the performances of both end-to-end image and video compression schemes have surpassed that of the advanced H.266/VVC under certain test conditions [10, 11, 26, 27].
For the evaluation of these image/video compression schemes in the practical application and standardization, they are usually benchmarked with objective and subjective metrics to evaluate their rate-distortion (RD) performance and a trade-off between coding efficiency and reconstructed quality. To sufficiently consider the effectiveness of these quality assessments, the test results of representative image/video test datasets are the key to reflecting the practicability and generalization of the researcher’s scheme.
In this paper, a new image/video dataset, named USTC-TD, is proposed for testing and evaluating the practical image/video coding algorithms. USTC-TD contains 40 images and 10 video sequences with a wide content coverage. For the image dataset, each image has a high spatial resolution (4K) in RGB, YUV444/420 color space, and PNG/YUV file format. For the video dataset, each video sequence consists of 96 frames, and each frame is captured at 30 frames per second (fps) with 1080p spatial resolution in RGB, YUV444/420 color space, and PNG/YUV file format. For the construction of datasets, the data is collected with the specific design of different content factors (environmental/imaging-related factors), which aims to cover as close as possible to the real-world coding transmission scenes. Compared with the previous common test image/video datasets [28, 29, 30, 31, 32, 33, 34], we use different quantitative criteria to comprehensively evaluate the diversity of the proposed dataset from the perspective of spatial, temporal, colorfulness, and lightness information, and demonstrate the wide coverage for various image/video features.
TABLE I: Common Test Datasets of Image Compression
Dataset Resolution Number Color Space Bit Depth Setting Characteristic
Kodak[28] 768×512 24 RGB 24 - Rich Texture Various Scenery Appropriate Exposure
Tecnick[29] 1200×1200 100 RGB 24 Sampling
CLIC[30] 1189×1789 (AVG) 41 RGB 24 Valid-Professional
TABLE II: Common Test Datasets of Video Compression
Dataset Resolution Number FPS Length Characteristic
UVG[31] 3840×2160 16 50/120 5s-12s Fast/Slow Motion Diverse Video Scenes Appropriate Exposure
MCL-JCV[32] 1920×1080 30 - 5s
HEVC CTC[33], VVC CTC[34] 240P-4K 41 24-60 150-600 Frames
In addition, we establish baselines and evaluate the classic standardized compression schemes [35, 1, 2, 3, 4, 5, 36] and recent learned image/video compression schemes [6, 7, 8, 9, 10, 11, 12, 16, 19, 20, 21, 22, 23, 25, 26, 27] under different metrics (PSNR, MS-SSIM[37]), and then benchmark and analyze their performance on the proposed dataset to shed light on the future research and development of image/video coding. The benchmark data and test scripts are online available with the proposed dataset and released on the open-sourced website for researchers to reproduce conveniently.
We hope the proposed test datasets allow researchers to make more well-informed decisions under efficient evaluation, and guide the innovation and improvement of future schemes and experiments. In summary, our contributions are as follows:
  • We build a new image/video compression test dataset (named USTC-TD), which focuses on the diversity of various content factors.
  • We conduct a comprehensive evaluation of the proposed dataset by using different quantitative criteria, and demonstrate the wide coverage of various image/video features.
  • We conduct a comprehensive evaluation of the advanced image/video compression schemes on the proposed dataset, and establish an extended baseline for the evaluated image/video coding schemes benchmarked on USTC-TD.
  • Taking a close look at USTC-TD, we analyze the benchmarked performance and shed light on the future research and development of image/video coding.
The remainder of the paper is structured as follows. Section II mentions the background of previous compression-related test datasets. Section III summarises the data collection process of the proposed dataset. Section IV introduces the construction of the image/video dataset of USTC-TD, and analyzes the characteristics of the proposed dataset. Section V presents the experimental configuration and the evaluation of the advanced compression schemes, and further analyzes their performance, limitation, and inspiration. Section VI concludes the paper and presents some suggestions for future work. Finally, we mention the copyright for the usage of USTC-TD.
II Background
In the past twenty years, with the rapid development of multimedia data over the advanced exhibition devices, resolutions, frame rates, dynamic range and viewpoints, the transmission/storage quantity of multimedia data is progressively accompanied by dramatic increases in the requirement of users. As the powerful multimedia data transmission/storage tool, lossy/lossless image/video compression has become the primary driver for reducing the internet bandwidth and storage. For the standardization activities and research of compression-related systems, image/video test (evaluation) dataset is a critical component for optimizing the performance and reflecting the practicability and generalization of different compression schemes. Here we review the image/video test dataset commonly used by standards and researchers in the past, and summarize their characteristics.
Image Compression Test Dataset. For the evaluation of previous image compression schemes, Kodak[28], Tecnick[29] (sampling setting), CLIC (professional setting) [30] are commonly used, the setting is mentioned in the Table I , and the introduction is summarised below:
  • Kodak[28] is a commonly used true color set of images released for various testing purposes and benchmarks, it contains 24 images with RGB format. The images are all photographic type and continuous tone. Many sites use them as a standard test suite for compression testing.
  • Tecnick[29] is a huge collection of sample images designed for quality assessment of different kinds of displays and image processing techniques. The sampling setting is widely used on testing resampling algorithms.
  • CLIC[29] is a high-quality image set collected from Unsplash[38], and contains images of similar quality from potentially different sources. It has been successfully applied in the workshop and challenge on learned image compression (CLIC) of IEEE/CVF Computer Vision and Pattern Recognition Conference (CVPR) and Data Compression Conference (DCC).
In these image datasets, characteristics mainly focus on the different resolutions with more scene types, most of the test images are captured by high-definition lens in specific scenes. These datasets aim to evaluate the basic ability of image compression algorithms to remove intra-frame redundancy for different scenarios, but the limited diversity of content factors makes it difficult to evaluate the robustness of algorithms.
Video Compression Test Dataset. For the evaluation of previous video compression schemes, UVG[31], MCL-JCV[32], HEVC Common Test Conditions (CTC)[33], VVC CTC [34] are commonly used, the settings are mentioned in the Table II , and the introduction is summarised below:
  • UVG[31] contains 16 test video sequences. They are captured with Sony F65 video camera in 16-bit F65RAW-HFR format and converted to YUV420 videos by ffmpeg tool[39]. It is widely used in the evaluation of advanced learned video compression methods[12, 16, 19, 20, 21, 22, 23, 25, 26, 27].
  • MCL-JCV[32] is a compressed video quality assessment dataset based on the just noticeable difference (JND) model. All provided video sequences are available to the public with measured raw JND data for each test subject and allow users to do their own processing.
  • HEVC CTC, and VVC CTC[33, 34] define the common test conditions and test sequences for the standardization activities of H.265/HEVC [1] and H.266/VVC [2], and protect the core experiments in a well-defined rule. It promotes the upgrading of many technologies in standardization, and has been widely used in compression-related systems.
In these video test datasets, the characteristics mainly focus on the various video contents, including simple/complex motion and poor/high capture quality. Most of the test videos can only evaluate the basic ability of video compression-related algorithms to remove inter-frame redundancy for different scenarios under different video coding configurations, like motion estimation (ME), motion compensation (MC), and rate allocation/control (RC) technologies in low-delay (LD) and random access (RA) configurations, but these video contents with the limited types of temporal correlation make it difficult to evaluate the robustness of temporal property-related algorithms in video-based compression applications.
TABLE III: Camera Captured Parameters of USTC-TD 2022
Nikon-D3200 Specifications
Sensor Type CMOS (Nikon DX format)
Sensor Size 23.2mm×$\times$15.4mm
Effective Pixels 24.7million
Largest Image Size 6016$\times$4000
TABLE IV: Camera Captured Parameters of USTC-TD 2023
Nikon-Z-fc Specifications
Sensor Type CMOS (Nikon DX format)
Sensor Size 23.5mm×$\times$15.7mm (APS-C)
Effective Pixels 20.9million
Largest Image Size 5568$\times$3712
III Data Collection
In this section, we introduce the hardware, format and collection configuration of dataset collection.
III-A Camera and Format Configuration
The images and video sequences are captured by using Nikon-D3200 and Nikon-Z-fc for USTC-TD 2022 and 2023 datasets, and the specific camera parameters are shown in Table III and Table IV . For the format of images and video sequences in the dataset, they are transcoded from Raw camera format (DNG, MOV) and then converted to RGB, YUV 4:4:4, YUV 4:2:0 color space/format by using the ffmpeg [39] tool and the specific conversion standard of color space (BT.601, BT.709 [26, 40]).
III-B Collection Configuration
To develop a comprehensive and diverse image/video dataset, we consider the various content factors of collected data, including the environmental factors (e.g. scene type, texture, motion, view), imaging factors (e.g. illumination, lens, shadow), which cover as close as possible to the real-world coding transmission scenes. For each factor, we categorize it into different types implicit in the construction of our dataset, the categories are shown in Table V . According to these conditions, we choose more than twenty different scene types (e.g. dormitory, library, river bank, institutes, parks, classroom, street, vehicles), and adjust different camera parameters to capture. For each image, we take ten shots of the same scene at the same time to select the better one. For each video, we record five minutes for each scene with the same range of joint sense to select the ninety-six frames.
TABLE V: Collection Configuration of USTC-TD 2022 and 2023
Collection Configuration
Element Category Example
Texture
Structural,
Natural,
Geometric
Scenery,
People,
Gridding
Motion
Complex,
Medium,
Tiny
Occlusion,
Walking,
Chatting
View
Upward Level,
Horizontal Level,
Overhead Level
Building,
People,
Close Shot
Illumination
Appropriate Exposure,
Underexposure,
Overexposure
Natural Light,
Dark Light,
High Light
Lens
Moving,
Fix
Camera Motion,
Surveillance
Shadow
Hard,
Soft,
Cast
Camera Flash Lamp,
Natural Illumination,
Building Occlusion
IV Dataset Construction and Analysis
In this section, we introduce the construction of our proposed USTC-TD image and video datasets, and further analyze them based on the comparison with previous common test image/video datasets under different quantitative criteria.
IV-A Construction of USTC-TD 2022 and 2023 Image Dataset
Based on the characteristics of previous image datasets[28, 29, 30], our proposed dataset aims to cover various scenarios. Considering the various content factors, we combine different environmental factors and imaging factors in the collection process. For the diversity of environmental factors, we consider the scene type, texture, and view factors. For the diversity of imaging factors, we consider the resolution, illumination, and shadow factors. In Fig.  1 , we show the all collected image data of USTD-TD 2022, and in Table VI , we show the specific configuration of each image, and make it convenient to the researchers’ scheme design for different application scenes. The collected image data and configuration of USTC-TD 2023 are also mentioned in Fig.  1  and Table VII . Based on USTD-TD 2022, USTC-TD 2023 considers more extreme factors in real-world scenes.
Figure 1: Illustration of the image dataset in USTC-TD 2022 and 2023.
TABLE VI: The Configuration of USTC-TD 2022 Image Dataset
Images Scene Type Resolutions Texture Illumination View Shadow
USTC-2022-01 Scenery 4096×2160 Geometric Appropriate Exposure Horizontal Level Soft
USTC-2022-02 Scenery 4096×2160 Structural Appropriate Exposure Upward Level Soft
USTC-2022-03 Scenery 4096×2160 Structural Underexposure Horizontal Level Hard
USTC-2022-04 Scenery 4096×2160 Structural Appropriate Exposure Horizontal Level Soft
USTC-2022-05 People 4096×2160 Natural Overexposure Horizontal Level Hard
USTC-2022-06 Scenery 4096×2160 Geometric Appropriate Exposure Upward Level Cast
USTC-2022-07 Building 4096×2160 Geometric Appropriate Exposure Upward Level Cast
USTC-2022-08 Scenery 4096×2160 Geometric Appropriate Exposure Upward Level Cast
USTC-2022-09 People, Building 4096×2160 Natural Underexposure Horizontal Level Hard
USTC-2022-10 Building 4096×2160 Geometric Appropriate Exposure Upward Level Soft
USTC-2022-11 People, Scenery 4096×2160 Natural Appropriate Exposure Horizontal Level Soft
USTC-2022-12 People 4096×2160 Natural Underexposure Horizontal Level Hard
USTC-2022-13 Scenery 4096×2160 Structural Appropriate Exposure Upward Level Cast
USTC-2022-14 People, Vehicle 4096×2160 Natural Appropriate Exposure Horizontal Level Cast
USTC-2022-15 People 4096×2160 Natural Underexposure Upward Level Hard
USTC-2022-16 People 4096×2160 Natural Appropriate Exposure Horizontal Level Soft
USTC-2022-17 People, Building 4096×2160 Natural Overexposure Horizontal Level Hard
USTC-2022-18 People, Building 4096×2160 Natural Appropriate Exposure Horizontal Level Soft
USTC-2022-19 People, Building 4096×2160 Natural Overexposure Horizontal Level Soft
USTC-2022-20 People, River 4096×2160 Natural Appropriate Exposure Overhead Level Cast
TABLE VII: The Configuration of USTC-TD 2023 Image Dataset
Images Scene Type Resolutions Texture Illumination View Shadow
USTC-2023-01 People, Room 3840×2160 Natural Appropriate Exposure Horizontal Level Hard
USTC-2023-02 Scenery 3840×2160 Geometric Underexposure Overhead Level Cast
USTC-2023-03 Scenery 3840×2160 Structural Underexposure Horizontal Level Soft
USTC-2023-04 Bicycle, Dense Objects 3840×2160 Geometric Appropriate Exposure Horizontal Level Soft
USTC-2023-05 Plant, Dense Textures 3840×2160 Structural Overexposure Horizontal Level Soft
USTC-2023-06 Water Wave 3840×2160 Geometric Appropriate Exposure Overhead Level Cast
USTC-2023-07 Building 3840×2160 Geometric Overexposure Upward Level Soft
USTC-2023-08 Plant, Dense Textures 3840×2160 Structural Appropriate Exposure Overhead Level Soft
USTC-2023-09 Scenery, People 3840×2160 Natural Underexposure Horizontal Level Hard
USTC-2023-10 People 3840×2160 Natural Overexposure Overhead Level Soft
USTC-2023-11 People 3840×2160 Natural Overexposure Horizontal Level Soft
USTC-2023-12 People 3840×2160 Natural Underexposure Upward Level Soft
USTC-2023-13 People 3840×2160 Natural Appropriate Exposure Horizontal Level Soft
USTC-2023-14 Building 3840×2160 Geometric Appropriate Exposure Horizontal Level Cast
USTC-2023-15 Plant 3840×2160 Structural Appropriate Exposure Horizontal Level Soft
USTC-2023-16 People, Occlusion 3840×2160 Natural Appropriate Exposure Horizontal Level Soft
USTC-2023-17 Close Shot 3840×2160 Natural Appropriate Exposure Horizontal Level Soft
USTC-2023-18 People 3840×2160 Natural Appropriate Exposure Horizontal Level Soft
USTC-2023-19 Close Shot 3840×2160 Natural Appropriate Exposure Horizontal Level Hard
USTC-2023-20 Close Shot 3840×2160 Natural Appropriate Exposure Overhead Level Cast
Figure 2: Illustration of the frame 01, 06, 11 of each video sequence in video test dataset of USTC-TD 2023.
TABLE VIII: The Configuration of USTC-TD 2023 Video Dataset
Video Sequences Color Space Motion Scene Types Resolutions Quality Texture View Lens
USTC-Badminton YUV420, 444, RGB Medium People, Sport 1920×1080 High Natural Horizontal Level Moving
USTC-BasketballDrill YUV420, 444, RGB Medium People, Sport 1920×1080 High Natural Horizontal Level Moving
USTC-BasketballPass YUV420, 444, RGB Medium People, Sport 1920×1080 High Natural Horizontal Level Moving
USTC-BicycleDriving YUV420, 444, RGB Complex People, Daily Life 1920×1080 High Natural Horizontal Level Moving
USTC-Dancing YUV420, 444, RGB Complex People, Sport 1920×1080 High Natural Horizontal Level Fix
USTC-ParkWalking YUV420, 444, RGB Complex People, Daily Life 1920×1080 High Natural Horizontal Level Moving
USTC-Running YUV420, 444, RGB Complex People, Sport 1920×1080 High Natural Horizontal Level Moving
USTC-ShankingHand YUV420, 444, RGB Complex People, Daliy life 1920×1080 High Natural Horizontal Level Moving
USTC-Snooker YUV420, 444, RGB Tiny Sport 1920×1080 High Natural Horizontal Level Moving
USTC-FourPeople YUV420, 444, RGB Tiny People 1920×1080 High Natural Horizontal Level Fix
Compared to the previous image datasets[28, 29, 30], more specific content factors are considered in our dataset. For example, in USTC-2022-09 and USTC-2022-05, we capture the low-light image with underexposure and high-light image with overexposure, which is a challenge for the generalization of many researchers’ image compression schemes [41, 42]. In USTC-2023-16, we capture the scenes with the object occlusion and the spatial-wise correlation becomes low, which is a challenge for traditional intra-prediction schemes [43] in the traditional codec[1, 2, 3, 4, 5]. We hope these specific testing sets can help the researchers discover the problem related to spatial characteristics in their image compression scheme.
IV-B Construction of USTC-TD 2023 Video Dataset
Based on the characteristics of previous video datasets[31, 32, 33, 34], our proposed dataset aims to cover more typical characteristics of video content. Compared to the image data, temporal-domain properties are unique to video, especially in the diverse motion types with more environmental and imaging factors in natural videos. There are usually multiple moving objects of arbitrary shapes and various motion types in video frames, leading to complex motion fields, which challenge the video coding schemes[44, 45, 46, 47]. Therefore, we simulate the video data with various temporal correlation types, including different kinds of motion types and lens motion.
In Fig. 2 , we show the partial frames of all collected video sequences in the USTD-TD 2023 dataset, and the specific configuration of each video in Table VIII , which make it convenient for the scheme design of research in real-world scenes under different temporal correlation types. Note that the different motion types are obviously classified in the Table VIII with different colors.
Different from the previous video dataset, we add more specific temporal correlation types in our proposed video dataset. For example, in USTC-BicycleDriving, we capture the video with a fast scene change (lens motion), high-speed moving objects, and object occlusion, which is a challenge for many inter-alignment schemes of submitted solutions in VCIP Challenge 2023 [48] and many optical flow-based video compression schemes [12, 49, 16, 19, 20, 21, 22, 23, 25, 26, 27] (the detailed performance analysis is mentioned in Section V.B). For the performance of different schemes of this sequence, the learned video compression schemes are far inferior to the traditional codecs [1, 2, 3, 4, 5]. In USTC-Snooker, we capture the scenes with the tiny motion and fast lens motion, which is also a challenge for optical flow-based schemes. At present, most of the learned video compression schemes use the optical flow-based alignment[50, 51], the optical flow-based motion estimation is difficult to capture the tiny motion and further influences their performance. Therefore, we put forward our proposed video dataset with the above specific designs, and hope the efficient testing datasets can help the researchers discover the problem related to temporal characteristics on their video compression scheme.
Figure 3: The visualization of the evaluation of spatial information (SI) and colorfulness (CF) features on different image test datasets. Scatter diagram represents the SI versus CF, and corresponding convex hulls indicates the coverage of different datasets. The histogram represents the number of images under different SI scores.
Figure 4: The visualization of the evaluation of lightness information (LI) and CF features on different image test datasets. Scatter diagram represents the LI versus CF, and corresponding convex hulls indicates the coverage of different datasets. The histogram represents the number of images under different LI scores.
Figure 5: The visualization of the evaluation of temporal information (TI) and SI features on different video test datasets. Scatter diagram represents the TI versus SI, and corresponding convex hulls indicates the coverage of different datasets. The histogram represents the number of videos under different TI scores.
IV-C Analysis of USTC-TD Dataset
In the above subsections, the construction of each dataset of USTC-TD is introduced, and we also point out some specific characteristics of the proposed dataset. To comprehensively verify the wide coverage of our proposed dataset for various content factors and qualitatively analyze the superiority of USTC-TD, we evaluate the USTC-TD on different image/video features and compare it with the previous image/video common test datasets (image datasets: Kodak[28], CLIC[30], Tecnick[29], video datasets: HEVC CTC[33], VVC CTC[34], MCL-JCV[32], UVG[31]). For analysis of image/video features, we select the spatial information (SI)[52], colorfulness (CF)[52], lightness information (LI)[53], and temporal information (TI)[31] to characterize each dataset along the dimensions of space, color, lightness, and temporal correlation, which are commonly used to evaluate the quality of dataset [54, 55, 31]. The definitions of the evaluation of these features can be found in [52, 31, 56, 53, 57], and the detailed calculation schemes are as follows:
  • Spatial information (SI): SI is used as a representation of edge energy[58]. Followed by [52], the SI is defined as the root mean square of edge magnitude over the luma component of an image or a video frame:
    $Score_{SI}=\sqrt{\frac{L}{1080}}\sqrt{\sum{\frac{S_{r}^{2}}{P}}},$ (1)
    where $S_{r}=\sqrt{S_{v}^{2}+S_{h}^{2}}$ indicates the edge magnitude at each pixel. $S_{v}$ and $S_{h}$ indicate the images/video frames filtered with vertical and horizontal Sobel kernels, respectively. $P$ indicates the total number of pixels in the filtered image, and $L$ indicates the vertical resolution. The normalization factor $\sqrt{\frac{L}{1080}}$ is used to reduce the scale and resolution dependence of SI. For video datasets, followed by [31], SI is taking the maximum of the results of all video frames.
  • Colorfulness (CF): CF is used as a representation of the variety and intensity of colors in the image. Followed by [59, 31], CF is defined as
    $Score_{CF}=\sqrt{\sigma_{rg}^{2}+\sigma_{by}^{2}}+0.3\sqrt{\mu_{rg}^{2}+\mu_{by}^{2}}$ (2)
    where opponent color spaces ($rg$, $by$) are defined in RGB color space. To be special, $rg=R-G$ and $by=0.5\,(R+G)-B$.
  • Lightness information (LI): LI is used as a representation of lightness variation. To measure the lightness information, we adopt the root mean square (RMS) contrast[57], the LI is defined as the standard deviation of the pixel intensities:
    $Score_{LI}=\sqrt{\frac{1}{MN}\sum_{i=0}^{N-1}\sum_{j=0}^{M-1}(I_{i,j}-\hat{I})^{2}},$ (3)
    where the intensities ($I_{i,j}$) are the $i$-th and $j$-th elements of the two-dimensional image of size $M$ by $N$. $\hat{I}$ is the average intensity of all pixel values in the image. The pixel intensities of the image ($I$) are normalized in the range $[0,1]$.
  • Temporal information (TI): TI is used as a representation of temporal variation. Followed by [56, 31], TI is defined as the maximum amount of temporal variation between successive frames $F_{n-1}$ and $F_{n}$:
    $Score_{TI}=\max_{1\leq n\leq N-1}\left\{\underset{\begin{subarray}{c}0\leq i\leq W-1\\ 0\leq j\leq H-1\end{subarray}}{\mathrm{std}}\left[F_{n}(i,j)-F_{n-1}(i,j)\right]\right\}$ (4)
    where $W$, $H$, $N$ denote the frame width, height and the number of total frames, respectively.
TABLE IX: Quantitative Results of the USTC-TD 2022 and 2023 Image Datasets. Note the Highest values are Represented in Red,
and the Lowest Values are Represented in Blue
USTC-TD 2022 USTC-TD 2023
Image SI CF LI Image SI CF LI
01 13.57 167.24 0.27 01 12.81 155.76 0.24
02 10.71 181.37 0.15 02 13.13 162.53 0.21
03 11.08 161.07 0.22 03 12.22 191.29 0.21
04 12.13 191.26 0.21 04 14.44 171.29 0.27
05 12.45 178.87 0.20 05 15.17 102.66 0.17
06 10.74 187.86 0.30 06 14.12 182.55 0.25
07 13.93 168.51 0.27 07 9.99 173.61 0.35
08 11.17 191.82 0.26 08 15.16 135.16 0.14
09 12.49 164.82 0.25 09 13.02 187.97 0.33
10 10.58 169.67 0.22 10 12.97 168.71 0.30
11 13.63 171.74 0.18 11 13.65 157.03 0.25
12 12.89 112.00 0.21 12 12.28 180.91 0.31
13 10.59 186.17 0.19 13 13.77 128.10 0.27
14 14.00 175.38 0.24 14 12.67 169.37 0.27
15 12.29 162.23 0.21 15 14.64 150.07 0.18
16 11.24 72.44 0.25 16 13.03 149.98 0.30
17 10.62 159.53 0.35 17 14.26 84.85 0.23
18 14.09 173.38 0.18 18 12.23 74.75 0.25
19 12.33 176.98 0.26 19 12.76 142.93 0.24
20 12.70 180.76 0.25 20 13.62 162.03 0.22
TABLE X: Quantitative Results of the USTC-TD 2023 Video Dataset.
Note the Highest values are Represented in Red, and
the Lowest Values are Represented in Blue
USTC-TD 2023
Video Sequence SI TI
USTC-Badminton 67.37 27.91
USTC-BasketballDrill 130.31 42.25
USTC-BasketballPass 94.63 46.92
USTC-BicycleDriving 38.66 52.70
USTC-Dancing 74.34 19.61
USTC-FourPeople 45.49 9.54
USTC-ParkWalking 72.48 40.22
USTC-Running 86.52 34.84
USTC-ShakingHands 100.67 44.11
USTC-Snooker 41.62 29.63
The quantitative evaluation scores of different datasets are shown in Fig. 5 ,  5 ,  5 . From the comparison with other datasets, we find that USTC-TD can collaborate with other datasets to handle a wide coverage of different image/video features, which verifies the diversity of the proposed dataset.
Specifically, for the evaluation of USTC-TD image dataset, the scores of SI, LI, CF cooperate to evaluate its spatial, colorfulness and lightness diversity. Compared to the other image datasets[28, 30, 29], it exhibits SI scores ranging from 9 to 16, and is distinguished from other image datasets, as shown in Fig. 5 . The proposed image dataset incorporates more spatial diversity within the wide range of colorfulness diversity. In Fig. 5 , the proposed image dataset also shows a wide coverage of LI scores, ranging from 0.10 to 0.40, which aligns with the range of other image datasets and demonstrates that the proposed image dataset also exhibits excellent generalization of lightness diversity. For the evaluation of the USTC-TD video dataset, the scores of TI and SI cooperate to evaluate its spatial and temporal diversity, as shown in Fig. 5 . Compared to the other video datasets[33, 34, 32, 31], the USTC-TD exhibits a wide range of temporal variation with generalized spatial variation, ranging from 5 to 55. It further compensates for the absence of the 40 to 55 (higher) range in the temporal variation of other video datasets, which enables a wide coverage of temporal diversity for the comprehensive assessment of video compression-related algorithms.
In addition, we present the detailed variety distribution of these different features of USTC-TD image/video datasets, as shown in Table IX  and Table X . The higher/lower scores of different evaluative features are represented in different colors. Benefiting from the specific design of various content factors (environmental/imaging-related factors) of each image/video, we can find that the typical features of different content factors are evenly distributed in the captured image/video to ensure the diversity. For example, first, as mentioned in Section IV.A and Table VI , VII , IX , the setting of illumination (overexposure/underexposure) of USTC-2022-17, and USTC-2023-05, USTC-2023-09 enables the highest/lowest score of LI evaluative features, which demonstrates its specific design of imaging-related factors. Second, as mentioned in Section IV.B, Table VIII  and Table X  , the setting of complex motion (high-speed/tiny-speed moving objects) and lens (moving/fix) of USTC-BicycleDriving and USTC-FourPeople enables the higher/lower score of TI evaluative features, which demonstrates its specific design of environmental factors. We hope these specific design can help the researchers analyze and discover the bottlenecks of their research. The above analysis results and related codes are also open-sourced on the website.
Figure 6: Overall rate-distortion (RD) curves of advanced image compression schemes on PSNR and MS-SSIM metrics, the results are evaluated on USTC-TD image dataset 2022 and 2023.
V Experiments
In this section, first, we present the experimental configurations employed for the evaluation of compression schemes. Second, we evaluate the classic standardized compression schemes and recent advanced image/video compression schemes on the proposed dataset under different metrics, and benchmark their performance on our proposed dataset. Third, we analyze the benchmarked performance, and further point out some limitations and inspirations among these advanced schemes to shed light on the future research.
V-A Experimental Settings
In this subsection, first, we present the experimental settings of the evaluative compression schemes, including the selection of advanced image/video compression schemes and the training/testing configurations of these methods. Second, we introduce the evaluation metrics for these methods on our proposed dataset.
Figure 7: Overall rate-distortion (RD) curves of advanced video compression schemes on PSNR and MS-SSIM metrics, the results are evaluated on USTC-TD video dataset 2023.
V-A1 Selection of Evaluative Methods
For advanced image compression schemes, we select classic standardized schemes and advanced learned schemes. For traditional codecs, we select BPG [35] and H.266/VVC [2]. For learned image compression schemes, we select the Factorized Model[6], Hyperprior Model[6], Autoregressive Model[7], Cheng2020[8], iWave++[9], ELIC[10], and MLIC++[11]. For learned compression standardized schemes, we select the high-profile model of IEEE 1857.11[36] (iWave++).
For video compression schemes, we also select classic standardized schemes and advanced learned schemes. For standardized codecs, we select the H.265/HEVC[1], and H.266/VVC[2]. For learned video compression schemes, we select the DVC_Pro[12, 49], CANF-VC[16], DCVC[19], TCM-VC[20], DCVC-HEM[21], OOFE[22], SDD[23], VNVC[25], DCVC-DC[26], and DCVC-FM[27]. The detailed introduction and test instructions of these methods are mentioned in Section I of supplementary material.
V-A2 Testing Configurations of Evaluative Traditional Image Compression Schemes
For testing, the officially released BPG software, VTM-17.0 (H.266/VVC reference software) are chosen. For BPG, the default configuration is used, and the internal color space is set to YUV444/420 for the testing of BPG and BPG444. For VTM-17.0, the encoder_intra_vtm configuration is used, and the internal color space is set to YUV444. For the different source formats of these testing datasets (USTC-TD and [28, 29, 30]), we convert them to the YUV444 color space for the input of VTM-17.0 by using the ffmpeg [39] tool (BT.601 conversion standard by default).
V-A3 Testing Configurations of Evaluative Traditional Video Compression Schemes
For testing, the officially released HM-16.20 (HEVC reference software) and VTM-13.2 are chosen. For the setting of HM-16.20, the encoder_lowdelay_main_rext configuration is used. For the setting of VTM-13.2, the encoder_lowdelay_vtm configuration is used. The internal color space is set to YUV444. For the different source formats of these testing datasets (USTC-TD and [31, 32, 33, 34]), we also convert them to the YUV444 color space as the input of the above traditional codecs by using the ffmpeg [39] tool.
Figure 8: Specific rate-distortion (RD) curves of advanced image compression schemes on partial evaluative images under PSNR metric.
Figure 9: Specific rate-distortion (RD) curves of advanced video compression schemes on partial evaluative videos under PSNR metric.
V-A4 Training and Testing Configurations of Evaluative Learned Image Compression Schemes
For training, these learned image compression models are optimized by mean squared error (MSE) or multi-scale structural similarity index measure (MS-SSIM), and the Flicker2W [60] dataset is used as the training dataset. These models are optimized by using the Adam Optimizer [61], with a batch size of 8 and a patch size of $256\times 256$. They are optimized for around 1.2 million iterations, starting with an initial learning rate of $10^{-4}$. The learning rate is reduced to $10^{-5}$ after 400 epochs and further down to $10^{-6}$ after 30 epochs. The setting of $\lambda$ is set to $\{0.001,0.004,0.024,0.080,0.200\}$ for iWave++, and $\{0.0018,0.0035,0.0067,0.0130,0.0250,0.0483\}$ for other schemes. For testing, the officially released model of MLIC++ and iWave++, and the reproduced model of ELIC are used. The model of other methods is provided by CompressAI [62].
V-A5 Training and Testing Configurations of Evaluative Learned Video Compression Schemes
For training, these learned video compression models are mainly optimized by MSE or MS-SSIM, and the Viemo-90k[63] is used as the training set. For the testing of USTC-TD, the officially released models of these schemes are used. For the testing of the variable-rate model, the bitrate points are aligned to that of the traditional codec. For the testing of other models, we directly use the released model. For the different source formats of these testing datasets, we convert them from YUV444 to the RGB / YUV420 color space as the input of different learned video codecs by using the BT.709 conversion standard (adopted in JPEG AI[40]) / ffmpeg [39] tool.
V-A6 Evaluation Metrics
We use PSNR and MS-SSIM[37] to measure the quality of the coded frames in the comparison to the original frames. Bits per pixel (bpp) is used to measure the number of bits for encoding each pixel in each image or video frame. The Bjontegaard Delta bitrate (BD-rate)[64] is used to compare the performance of different compression schemes, where negative numbers indicate bitrate saving and positive numbers indicate bitrate increasing. For the evaluation of image and video compression schemes, the PSNR and MS-SSIM are both calculated and compared in RGB color space. For image, the conversion process of different color spaces is performed by using the ffmpeg tool [39]. For video, the conversion process is aligned to the setting mentioned in DCVC-DC[26] by using the BT.709 conversion standard.
V-B Experimental Results
In this subsection, we establish the baselines and benchmark the performance of advanced image/video compression schemes on USTC-TD image/video datasets, and further analyze their performance.
TABLE XI: BD-rate (%) Comparison for PSNR. The anchor is VTM.
Dataset VTM HM DVC_Pro DCVC CANF-VC TCM-VC VNVC DCVC-HEM OOFE SDD DCVC-DC DCVC-FM
HEVC Class B 0.0 39.0 188.6 115.7 58.2 32.8 24.6 -0.7 -15.3 -13.7 -13.9 -8.8
HEVC Class C 0.0 37.6 202.8 150.8 73.0 62.1 48.2 16.1 -2.2 -2.3 -8.8 -5.0
HEVC Class D 0.0 34.7 160.3 106.4 48.8 29.0 22.4 -7.1 -23.0 -24.9 -27.7 -23.3
HEVC Class E 0.0 48.6 429.5 257.5 116.8 75.8 66.8 20.9 -0.4 -8.4 -19.1 -20.8
HEVC Class RGB 0.0 44.0 186.8 118.6 87.5 25.4 16.0 -15.6 -17.5 -17.5 -27.9 -18.6
UVG 0.0 36.4 218.7 129.5 56.3 23.1 18.0 -17.2 -22.3 -19.7 -25.9 -20.5
MCL-JCV 0.0 41.9 163.6 103.9 60.5 38.2 30.2 -1.6 -5.8 -7.1 -14.4 -7.4
USTC-TD 0.0 46.9 286.7 133.0 69.3 69.8 63.9 17.0 9.5 4.5 7.1 23.1
Average 0.0 41.1 229.6 139.4 71.3 44.5 36.3 1.5 -9.6 -11.1 -16.3 -10.2
TABLE XII: BD-rate (%) Comparison for MS-SSIM. The anchor is VTM.
Dataset VTM HM DVC_Pro DCVC CANF-VC TCM-VC VNVC DCVC-HEM OOFE SDD DCVC-DC DCVC-FM
HEVC Class B 0.0 36.8 67.0 35.9 25.5 -20.5 -33.1 -47.4 -15.1 -48.0 -53.0 -12.5
HEVC Class C 0.0 38.7 61.1 24.9 17.7 -21.7 -29.3 -43.3 -15.9 -49.6 -54.6 -18.0
HEVC Class D 0.0 34.9 25.3 2.7 1.5 -36.2 -41.1 -55.5 -28.6 -60.0 -63.4 -30.6
HEVC Class E 0.0 38.4 195.8 90.0 114.9 -20.5 -0.4 -52.4 -9.3 -51.5 -60.7 -32.6
HEVC Class RGB 0.0 37.3 66.8 43.7 52.9 -21.1 -32.4 -45.8 -16.7 -46.3 -54.4 -16.6
UVG 0.0 37.1 74.6 11.9 33.1 -6.0 -15.2 -32.7 -10.6 -34.2 -36.7 -7.3
MCL-JCV 0.0 43.7 46.1 39.1 11.7 -18.6 -29.0 -44.0 -2.5 -46.3 -49.1 -5.0
USTC-TD 0.0 47.9 38.6 -18.1 77.5 -41.1 -62.1 -62.3 8.4 -66.3 -68.6 37.1
Average 0.0 39.4 71.9 28.8 41.9 -23.2 -30.3 -47.9 -11.3 -50.3 -55.1 -10.7
V-B1 Evaluation and Analysis of Advanced Image Compression Schemes on USTC-TD Image Datasets
Taking bpp as the horizontal axis and the reconstructed PSNR/MS-SSIM as the vertical axis, we present the rate and distortion curves of different image compression schemes over USTC-TD 2022 and 2023 image datasets in Fig. 6 . From the overall results, we can find that the partial learned schemes (iWave++[9], ELIC[10], MLIC++[11]) can outperform the traditional image compression schemes and achieve better compression performance than H.266/VVC on proposed datasets under different metrics, which show its powerful potential. Here we further analyze their performance from the perspective of content factors of different test images of USTC-TD.
As shown in Fig. 8 , the detailed RD curves of some test images (USTC-2022-12, USTC-2023-15, USTC-2022-17, USTC-2023-07) with some special phenomenons are illustrated, and the results of each test image are presented in the Section II of supplementary material. Based on the performance comparison of these schemes and feature analysis of proposed datasets (Section IV.C), the conclusions mainly include the following four aspects:
  • (1) The learned schemes show the good potential on some complex scenarios. For example, as shown in the Fig. 8  (a) and (b), compared to the overall results (Fig. 6 ), more learned schemes (Autoregressive Model[7], Cheng2020[8], iWave++[9], ELIC[10], MLIC++[11]) perform better than the traditional schemes on some test images with specific features of environment-related factors (Table IX ), such as the USTC-2022-12 with the lower scores of CF, the USTC-2023-15 with the higher scores of SI. Meanwhile, the results also demonstrate the previous image train/test datasets [28, 29, 30] can guide the researcher to train/evaluate the basic ability of intra-frame redundancy removal of their schemes to some extent.
  • (2) The traditional schemes show the powerful generalization ability in some extreme scenarios. For example, as shown in the Fig. 8  (c) and (d), compared to the overall results (Fig. 6 ), the generalization ability of learned schemes is lacking to handle the evaluative images with extreme mixture features of environment/imaging-related factors well (Table IX ), such as the USTC-2022-17 with the lower scores of SI, CF and the higher scores of LI, the USTC-2023-07 with the lower scores of SI and the higher scores of LI. Although the performance of learned schemes surpasses the traditional schemes in general, they are still limited to some extreme scenarios.
  • (3) Based on the analysis of different features (SI, CF, LI, TI) of proposed datasets (Section IV.C), the detailed characteristics of different test images can efficiently assist the researcher to analyze the detailed bottleneck of their compression scheme.
  • (4) Compared to the performance of these schemes on different datasets[28, 29, 30], the proposed datasets can collaborate with other datasets to handle a wide coverage of performance evaluation, which also demonstrates the efficiency of specific design of the proposed datasets’ different factors/features.
V-B2 Evaluation and Analysis of Advanced Video Compression Schemes on USTC-TD Video Dataset
Taking bpp as the horizontal axis and the reconstructed PSNR/MS-SSIM as the vertical axis, we present the rate and distortion curves and BD-rate results of different video compression schemes over the USTC-TD 2023 video dataset in Fig. 7 , Table XI  and Table XII . From the overall results of PSNR metric, we can find that the performance of the advanced traditional schemes is better than that of the all advanced learned schemes on the proposed dataset, which is opposed to the performance on other datasets [31, 32, 33, 34] . From the overall results of MS-SSIM metric, the performance of the traditional schemes is lower than that of advanced learned schemes. Here we further analyze their performance from the perspective of typical characteristics of different test video contents of USTC-TD video dataset.
As shown in Fig. 9 , the detailed RD curves of some test videos (USTC-FourPeople, USTC-BasketballPass, USTC-Snooker, USTC-BicycleDriving) with some special phenomenons are illustrated, the results of each test video are also presented in the Section II of supplementary material. Based on the performance comparison of advanced schemes and feature analysis of proposed datasets (Section IV.C), the conclusions mainly include the following three aspects:
Figure 10: Visualization of limitation of partial advanced image compression schemes (ELIC[10], MLIC++[11]) on USTC-2022-03 and USTC-2022-12.
Figure 11: Visualization of limitation of partial advanced video compression schemes (TCM-VC[20], DCVC-DC[26], DCVC-FM[27]) on USTC-BicycleDriving and USTC-Snooker.
  • (1) The traditional schemes show the robust generalization ability on various real-world scenarios. As shown in Table XI , different from the performance of advanced compression schemes on other datasets [31, 32, 33, 34] , the traditional schemes still achieve the state-of-the-art performance of these schemes on USTC-TD. Different from the other datasets, USTC-TD video dataset focuses on various temporal correlation types. Combined the attributes of different test videos of USTC-TD (Fig. 5 , Table VIII ), we find that the traditional schemes can handle more scenarios with the various temporal features of motion-related elements (motion type, lens motion), such as the USTC-BicycleDriving with higher scores of TI and USTC-Snooker with specific design of lens motion (mentioned in Section IV.C). The performance of these test videos is shown in Fig. 9  (c) and (d), and the results demonstrate that the traditional schemes can handle these scenes with complex motion types robustly.
  • (2) The learned schemes show the optimistic potential on some scenarios with complex motion, such as the USTC-FourPeople with lower scores of TI, and the USTC-BasketballPass with higher scores of TI. These scenarios commonly appear in the previous test datasets[31, 32, 33, 34] , such as the FourPeople and BasketballPass in HEVC/VVC CTC, which can guide the design and the optimized target of deep network to handle these motion situations. The results of these specific scenarios demonstrate the basic evaluative ability of the previous datasets and the performance potential of the deep learning-based manner.
  • (3) Based on the analysis of video-related features (Section IV.C), our proposed dataset can make up more typical motion/temporal correlation-related real-world factors with other datasets to handle a wide coverage of performance evaluation, which also demonstrates the efficiency of the specific design of the proposed datasets’ different factors/features. Furthermore, compared to the performance of other datasets[31, 32, 33, 34] of different schemes mentioned in [7, 8, 9, 10, 11], as shown in Table XI  and Table XII , the different phenomenon between the performance of our proposed dataset and other datasets also verifies the efficiency of the proposed dataset.
V-C Limitation and Inspiration
In this subsection, based on the analysis of experimental results, we further analyze the limitations of evaluative image/video compression schemes, and point out some limitations and inspirations among these advanced compression schemes.
V-C1 Limitation and Inspiration of Image Compression
Based on the above analysis in Section V.B, we can find that the generalization ability of the learned codec is a major challenge for practical usage. Most existing methods focus on improving the compression performance while neglecting its generalization for various scenarios. As the situations mentioned in the item (1) and (2) of conclusions (Section V.B (1)), the generalization ability is challenged with the effective extension of the evaluation data. These problems mainly arise from the incompletion of training data and the constraint optimization direction of deep learning-based manner with the limited evaluation data. Here we further explore these problems based on the coding process of different compression schemes.
For example, as shown in Fig. 6, the performance of one bpp point of MLIC++ is lower than that of other schemes, but the performance of other bpp points is better. In detail, we further illustrate the detailed rate-distortion curves of each image, such as the RD curves of USTC-2022-12, USTC-2022-17, USTC-2023-07 shown in Fig. 8 (a), (b), and (c), we can find that one bpp model of MLIC++ all performs poorly on several specific images. To explore it, we visualize the bits allocation map of ELIC and MLIC++ on some test images, such as the cases of USTC-2022-03 and USTC-2022-12 shown in Fig. 10 . Compared to ELIC, MLIC++ allocates more bits to some flat areas, whereas these areas could be encoded with fewer bits. Inspired by them, the controllable model optimization, domain adaptation, and precise rate allocation of the learned compression models need to be further improved for future practical usage.
V-C2 Limitation and Inspiration of Video Compression
Based on the above analysis in Section V.B, as shown in Fig. 7 and Table XI , the performance of all learned video codecs is lower than that of traditional video codecs on USTC-TD, which is different from the performance phenomenon on other datasets. The reason mainly comes from that the learned video codecs perform poorly on some sequences with complex motion features. Here we further explore these problems based on the motion-related modules of different schemes.
As mentioned in Fig. 9  (c) and (d), the performance of the state-of-the-art learned video codecs is even lower than that of the H.265/HEVC[1]. Based on these observations, we visualize the video reconstructed frames of these video codecs as shown in Fig. 11 . From the comparison of different scenarios, we can find that the specific design of motion-related features (high-speed moving objects, object occlusion, and camera motion) bring severe motion blur in the temporal domain, which challenges the optical flow-based motion estimation/compensation module of learned video codecs that is difficult to estimate accurate motion vector prediction. Therefore, we further illustrate the estimated motion vectors of these traditional and learned codecs in Fig. 11 , it obviously observes that the motion field of learned video compression schemes performs wrong and disordered, which further demonstrates that the flow-based motion-related modules of learned video codecs are difficult to handle the complex motion situations.
TABLE XIII: BD-rate (%) Results of Different Training Strategies of Optical Flow-related Module of DCVC-DC
on USTC-TD [26] under PSNR Metric
      Scheme       BD rate (%)
      DCVC-DC       7.1%
      DCVC-DC + Flow Pre-training       -1.6%
      $\triangle$       -8.7%
To further verify the performance impact of these problems, we tentatively design the experiment to optimize the optical flow-related module of different learned video codecs. We set the state-of-the-art scheme (DCVC-DC [26]) as the anchor, and use the motion vectors of H.266/VVC as the optimized target of the optical flow-based motion estimation module (Spynet[51]) in the offline pre-training stage, instead of the usage of EPE loss for the training of these optical flow-based modules. The performance is shown in Table XIII . Inspired by the results, it verifies that the motion modeling and training strategy of learned video compression models are necessary to be further improved for practical usage in the future.
VI Conclusion
In this paper, we propose a test dataset (named USTC-TD) for compression-related research, which covers more diverse content factors. To evaluate the efficiency of USTC-TD, we qualitatively evaluate the USTC-TD on different image/video features and compare it with the previous image/video common test datasets to verify its wide coverage. In addition, we evaluate the advanced compression methods under different metrics benchmarked on USTC-TD, and further analyze their performance to point out the inspirations for future compression-related research.
In the present dataset construction process, we only consider the basic image and video test datasets. In the future, we plan to progressively extend the annotation datasets of USTC-TD for image/video coding for machine (ICM/VCM)[65, 66, 67], such as object segmentation [68], object detection [69], action recognition [70], et al, and the reconstruction dataset of video enhancement, such as image/video super-resolution[71, 72], denoising[73], warping[74], et al, for the testing of compression-related downstream researches, and provide a comprehensive baseline to promote the development of compression-related diverse tasks.
VII Copyright
The released images and sequences are captured and processed by the University of Science and Technology of China (USTC). All intellectual property rights remain with USTC.
The following uses are allowed for the contributed dataset: (1) Data (images and videos) may be published in research papers, technical reports, and development events. (2) Data (images and videos) may be utilized for standardization activities (e.g., ITU, MPEG, AVS, VQEG).
The following uses are NOT allowed for the contributed dataset: (1) Do not publish snapshots in product brochures. (2) Do not use video for marketing purposes. (3) Redistribution is not permitted. (4) Do not use it in television shows, commercials, or movies.